Problem 10 [Objective 1 potential energy of tis system Answers (a) -1.3 nJ (b)-43J 13a 1.3 nJ (d) 4.3 nJ (e) Zero 9 of 18
Answer #1

The total Potential energy is given by

U_{total} = \frac{1}{4\pi \epsilon_o \kappa} [\frac{Q_AQ_B}{r_{AB}}+\frac{Q_BQ_C}{r_{BC}}+\frac{Q_AQ_C}{r_{AC}}]

U_{total} = \frac{9*10^9}{4} [\frac{6*10^{-9}*(-4)*10^{-9}}{8}+\frac{(-4)*10^{-9}*(6)*10^{-9}}{\sqrt{8^2+6^2}}+\frac{6*10^{-9}*6*10^{-9}}{6}]

\Rightarrow U_{total} = 1.35*10^{-9} J =1.35 nJ

Answer: Option c)